Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, November 30, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 30, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/30/98 (Monday)

Segment 1 report for 11/30/98: The bore hole at Cuddy Creek Tributary (EMP 25.43) has been completed. On Tejon Bench excavation continued at EMP 18.09. Backfill continued to EMP 17.06. Welding, x-ray and coating occurred from EMP 16.57 to 16.54. On Grapevine Rd., coating occurred to EMP 15.95, welding to 15.83. Trench was excavated from the mainline to the tie-in point at Grapevine Pump Station (EMP 15.57). At the pump station, electrical work and pipe-fitting continued. The bore pit on the north side of Stream Crossing #9 (EMP 15.35) was completed. Tie-in of the pipe through the bore at Tecuya Creek/I-5 bore (EMP 9.85) began. On Hwy. 166, recountering of the ROW occurred from EMP .53 to .62 and pipe to be pushed through the bore was welded, x-rayed and coated. At Emidio Pump Station, work on the office interior continued.

Segment 2B (Castaic to Sylmar): Report for 11-30-98. Drilling crews at the Santa Clara river began reaming the bore with a 12" bit completing 500' today. In addition, crews worked to pump out drilling mud and install silt fence near potential spill sites. At Feedmill Rd, crews backfilled and asphalted across the roadway and continued a general clean up of the extra work space. At Calgrove Blvd, crews installed 200' of fiber optic duct, backfilled then asphalted 100' of the ditch. Trenching crews completed 300' to EMP 71.8 south of Weldon Cyn, pipe was laid in the ditch and welded to EMP 72.05, X-ray to EMP 72.14 and coating to EMP 72.23. Further, 400' of fiber optic duct and slurry were completed to near EMP 72.5.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 30 November 1998: 3A: Backfill crews moved from south Tujunga Wash to work on backfilling at Hollywood Way Bridge crossing. This crew also worked on closing up the north side of the San Fernando Bridge crossing, though the main focus was on Hollywood Way. Fiber optic crews pulled cable (through the black conduit) in Glendale, between Highland and Broadway. The tie in on the north side of Tujunga Wash was completed. Welding, x-ray, and coating worked at Pacoima Wash Bridge crossing. The bore at Tuxford Street was nearly completed today. 3B: Electrical work was done at the block valve at the end of Segment 3, MP 98.48. Saw cutting in the road to hook up to an electrical source was also done. 3C: Excavation crews began digging between the skip north of Hubbard and the bore site today. Welding, x-ray, and coating worked above the ditch on the pipe for this area. Backfill continued to work just south of the aqueduct crossing, as well as on clean up and grading. Welding, x-ray, and coating completed their work on the tie in to Segment 2. Work on the cover for the high speed channel continued. Two projects memos were issued today. The first addressed the need to contain and properly dispose and drips, leaks, and spills. The second was a reminder of the need for erosion control and filtration during dewatering operations.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page